OpenStudy (anonymous):

solve for sqrt of u^15

OpenStudy (anonymous):

\(15\div 2=?\)

OpenStudy (anonymous):

\[\sqrt{u ^{15}}\]

OpenStudy (solomonzelman):

question incomplete (to me at least).

OpenStudy (anonymous):

simplify

OpenStudy (solomonzelman):

You want to convert \(\large\color{slate}{ \sqrt{u^{15}} }\) to a fractional exponent?

OpenStudy (solomonzelman):

\(\large\color{slate}{ \sqrt[C]{A^{B}}=A^{B/C} }\)

OpenStudy (solomonzelman):

\(\large\color{slate}{ \sqrt[2]{u^{15}} =? }\)